# An Improved Privacy and Utility Analysis of Differentially Private SGD with Bounded Domain and Smooth Losses

### v2

@@ A.3 Lemmas for Utility Analysis
Lemma A.13.
(Implications of smoothness (Beck, 2017)). Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex function, differentiable over $\mathbb{R}^{d}$. Then, the following claims are equivalent:
  • $f$ is $L$-smooth.
@@
Lemma A.14.
(Implications of strong convexity (Beck, 2017)). Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a proper closed and convex function, differentiable over $\mathbb{R}^{d}$. Then, for a given $\mu>0$, the following claims are equivalent:
  • $f$ is $\mu$-strongly convex.
@@ Appendix B Proofs for Privacy Analysis
(1) Based on Lemma A.7, for the shifted Rényi divergence $\mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})$, there exist jointly distributed random variables ($\boldsymbol{\mu},\boldsymbol{\mu}^{\prime})$ such that $\operatorname{Pr}[||\boldsymbol{\mu}-\boldsymbol{\mu}^{\prime}||\leq z]=1$ and $\mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})=\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\boldsymbol{\nu}})$. Then, one may obtain
@@
$\mathcal{D}_{\alpha}^{((1+\eta L)z)}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})$ $\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})$ (38)
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime})}||\mathbb{P}_{\psi^{\prime\prime}(\boldsymbol{\nu})})$ $\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime}),\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\psi^{\prime\prime}(\boldsymbol{\nu}),\boldsymbol{\nu}})$ (39)
@@
$\mathcal{D}_{\alpha}^{((1+\eta L)z)}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})$ $\leq(1-q)\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}+q\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\boldsymbol{\nu}})$ (40)
@@
(2) Assuming $b\leq n/5$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\text{DP}}}{2C})$, and $\sigma_{\text{DP}}>\frac{8C}{b\sqrt{\beta}}$, the proof is similar to the previous case. Starting with Equation 38, we obtain
$\mathcal{D}_{\alpha}^{((1+\eta L)z)}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})$ $\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime}),\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu}),\boldsymbol{\nu}})$ (41)
@@
Then, consider real sequence $\{a_{t}\}_{i=0}^{T-1}$ such that $z_{t}=\sum_{i=0}^{t-1}(1+\eta L)^{t-i-1}(-a_{i})$ is non-negative for all $t$ and $z_{T}=0$. By this way, we have $z_{0}=0$ and $z_{t+1}=(1+\eta L)z_{t}-a_{t}$. The proof proceeds by induction, leveraging Lemma 2.6 and Lemma 3.2. Specifically, we have
$\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\boldsymbol{\theta}_{t+1}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{t+1}})$ $=\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\psi(\boldsymbol{\theta}_{t})+\boldsymbol{\varsigma}_{t}}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\theta}^{\prime}_{t})+\boldsymbol{\varsigma}_{t}})$ (43)
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})$ $\leq\mathcal{D}_{\alpha}^{(z_{0})}(\mathbb{P}_{\boldsymbol{\theta}_{0}}||\mathbb{P}_{\boldsymbol{\theta}_{0}^{\prime}})+\sum_{t=0}^{T-1}\frac{\alpha a_{t}^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}+\sum_{t=0}^{T-1}\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}$ (44)
@@
$\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\boldsymbol{\theta}_{t+1}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{t+1}})\leq\mathcal{D}^{(z_{t})}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{t}}||\mathbb{P}_{\boldsymbol{\theta}_{t}^{\prime}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}+\frac{\alpha a_{t}^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}.$ (46)
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})$ $\leq\mathcal{D}_{\alpha}^{(z_{\tau})}(\mathbb{P}_{\boldsymbol{\theta}_{\tau}}||\mathbb{P}_{\boldsymbol{\theta}_{\tau}^{\prime}})+\sum_{t=\tau}^{T-1}\frac{\alpha a_{t}^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}+\sum_{t=\tau}^{T-1}\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}$ (47)
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})$ $\leq\min_{\tau\in\{0,\cdots,T-1\}}\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}(T-\tau)+\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}$ (48)
@@ Appendix C Proofs for Utility Analysis
where $g(\boldsymbol{\theta}_{t})=\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}g_{\xi}(\boldsymbol{\theta}_{t})=\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}_{C}\left(\nabla l_{\xi}(\boldsymbol{\theta}_{t})\right)$ and $\boldsymbol{\zeta}_{t}\sim\mathcal{N}(\boldsymbol{0},\eta^{2}\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ is the Gaussian perturbation.
@@
$\mathbb{E}_{\boldsymbol{\zeta}_{t}}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ $\leq\mathbb{E}_{\boldsymbol{\zeta}_{t}}\left[\|\boldsymbol{\theta}_{t}-\eta g(\boldsymbol{\theta}_{t})+\boldsymbol{\zeta}_{t}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (50)
@@
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})+\eta^{2}C^{2}+d\eta^{2}\sigma_{\text{DP}}^{2},$
@@
$-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$ $=-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma_{\xi}\nabla l(\boldsymbol{\theta}_{t})+\gamma_{\xi}\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$ (51)
$\leq-2\eta\gamma_{\xi}[\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma_{\xi}\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$\leq-\frac{7\eta C}{4\|\nabla l(\boldsymbol{\theta}_{t})\|}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]+2\eta\gamma_{\xi}\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|$
$\leq-\frac{7\eta C}{4\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}+\frac{2\eta C\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{4}}}{7\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{3}{4}}-\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{4}}}\sqrt{\frac{2}{\mu}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
$\leq-\frac{7\eta C}{4\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}+\frac{4\eta C}{\sqrt{2}(7\sqrt{L}-\sqrt{\mu})}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
@@
where the third step is by the convexity and Cauchy–Schwarz inequality, and the fourth step is by Lemma A.13 and Lemma A.14.
@@
$-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\leq 2\eta C\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|\leq\frac{4\eta C}{\sqrt{2\mu}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})},$ (52)
by using Cauchy–Schwarz inequality and Lemma A.14. Here, we define $\kappa_{\xi}=\mathbbm{1}\{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|>5\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{4}}\}$. According to Lemma A.1, we have
$\operatorname{Pr}(\kappa_{\xi}=1)$ $=\operatorname{Pr}(\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|^{2}>25\sigma_{\text{SGD}}^{2}\sqrt{L/\mu})\leq\frac{\sigma_{\text{SGD}}^{2}}{25\sigma_{\text{SGD}}^{2}\sqrt{L/\mu}}=\frac{1}{25}\sqrt{\frac{\mu}{L}},$ (53)
@@
$\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ $\leq-\frac{13\eta C}{12\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\cdot\frac{24}{25}+\frac{4\eta C}{\sqrt{2\mu}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\cdot\frac{1}{25}\sqrt{\frac{\mu}{L}}$ (54)

### v3

@@ A.4 Lemmas for Utility Analysis
Lemma A.14 (Implications of smoothness (beck2017first)).
Let $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex function, differentiable over $\mathbb{R}^{d}$. Then, the following claims are equivalent:
  • $f(\cdot)$ is $L$-smooth.
@@
Lemma A.15 (Implications of strong convexity (beck2017first)).
Let $f(\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a proper closed and convex function, differentiable over $\mathbb{R}^{d}$. Then, for a given $\mu>0$, the following claims are equivalent:
  • $f(\cdot)$ is $\mu$-strongly convex.
@@ Appendix B Proofs for Privacy Analysis
Lemma B.1 (Full statement of Lemma 3.2).
Let $\psi(\cdot)$ and $\psi^{\prime}(\cdot)$ be two noisy update functions (c.f. (9)) of DPSGD based on adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, and $n$ be the size of the dataset $\mathcal{D}$. If the loss function is $L$-smooth (Assumption 3.1), for any random variables $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$, we have
$\mathcal{D}_{\alpha}^{((1+\eta L)z)}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})\leq\ \mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\textup{DP}}^{2}}.$ (34)
If we further assume the mini-batch size $b\leq\frac{n}{5}$, the RDP parameter $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and the per-dimension Gaussian noise scale $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$, then
$\mathcal{D}_{\alpha}^{((1+\eta L)z)}(\mathbb{P}_{\psi(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})\leq\ \mathcal{D}_{\alpha}^{(z)}(\mathbb{P}_{\boldsymbol{\mu}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{8\alpha C^{2}}{\beta n^{2}\sigma_{\textup{DP}}^{2}}.$ (35)
@@
[…]
@@
$\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime})}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu})})$
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime})}||\mathbb{P}_{\psi^{\prime\prime}(\boldsymbol{\nu})})$ (38)
$\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime}),\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\psi^{\prime\prime}(\boldsymbol{\nu}),\boldsymbol{\nu}})$
@@
$\leq(1-q)\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\boldsymbol{\nu}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}+q\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\boldsymbol{\nu}})$
@@
(2) Assuming $b\leq n/5$, $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\text{DP}}}{2C})$, and $\sigma_{\text{DP}}>\frac{8C}{b\sqrt{\beta}}$, the proof is similar to the previous case. Starting with (37), we obtain
$\leq\mathcal{D}_{\alpha}(\mathbb{P}_{\psi(\boldsymbol{\mu}^{\prime}),\boldsymbol{\mu}^{\prime}}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\nu}),\boldsymbol{\nu}})$
@@
Theorem B.2 (Full statement of Theorem 3.3).
Given the number of total SGD iterations $T$, dataset size $n$, batch size $b$, stepsize $\eta$, $\alpha>1$, gradient clipping threshold $C$, and noise scale $\sigma_{\textup{DP}}$, if the loss function is $L$-smooth (Assumption 3.1), then the DPSGD-GC algorithm satisfies $(\alpha,\varepsilon)$-RDP for
$\varepsilon=\mathcal{O}\left(\frac{\alpha C^{2}}{nb\sigma_{\textup{DP}}^{2}}T\right).$ (41)
If we further assume the mini-batch size $b\leq\frac{n}{5}$, RDP parameter $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and the per-dimension Gaussian noise scale $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$ with a constant $\beta\in(0,1)$, then
$\varepsilon=\mathcal{O}\left(\frac{\alpha C^{2}}{n^{2}\sigma_{\textup{DP}}^{2}}T\right).$ (42)
@@
Then, consider real sequence $\{a_{t}\}_{i=0}^{T-1}$ such that $z_{t}=\sum_{i=0}^{t-1}(1+\eta L)^{t-i-1}(-a_{i})$ is non-negative for all $t$ and $z_{T}=0$. By this way, we have $z_{0}=0$ and $z_{t+1}=(1+\eta L)z_{t}-a_{t}$. The proof proceeds by induction, leveraging Lemma 2.6 and 3.2. Specifically, we have
$\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\boldsymbol{\theta}_{t+1}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{t+1}})$ (44)
$=\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\psi(\boldsymbol{\theta}_{t})+\boldsymbol{\varsigma}_{t}}||\mathbb{P}_{\psi^{\prime}(\boldsymbol{\theta}^{\prime}_{t})+\boldsymbol{\varsigma}_{t}})$
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})$ (45)
$\leq\mathcal{D}_{\alpha}^{(z_{0})}(\mathbb{P}_{\boldsymbol{\theta}_{0}}||\mathbb{P}_{\boldsymbol{\theta}_{0}^{\prime}})+\sum_{t=0}^{T-1}\frac{\alpha a_{t}^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}+\sum_{t=0}^{T-1}\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}$
@@
Theorem B.3 ((Full statement of Theorem 3.4).
Given the number of total SGD iterations $T$, dataset size $n$, batch size $b$, stepsize $\eta$, $\alpha>1$, gradient clipping threshold $C$, diameter of the model-parameter domain $D$, and noise scale $\sigma_{\textup{DP}}$, if the loss function is $L$-smooth (Assumption 3.1), then the DPSGD-DC algorithm satisfies $(\alpha,\varepsilon)$-RDP for
$\varepsilon=\mathcal{O}\left(\frac{\alpha C^{2}}{nb\sigma_{\textup{DP}}^{2}}\min\left\{T,\frac{(1+\eta L)^{2}nbD^{2}}{\eta^{2}C^{2}}\right\}\right).$ (47)
If we further assume the mini-batch size $b\leq\frac{n}{5}$, RDP parameter $\alpha\leq\alpha^{*}(\frac{b}{n},\frac{b\sqrt{\beta}\sigma_{\textup{DP}}}{2C})$, and the per-dimension Gaussian noise scale $\sigma_{\textup{DP}}>\frac{8C}{b\sqrt{\beta}}$ with a constant $\beta\in(0,1)$, then
$\varepsilon=\mathcal{O}\left(\frac{\alpha C^{2}}{n^{2}\sigma_{\textup{DP}}^{2}}\min\left\{T,\frac{(1+\eta L)^{2}n^{2}D^{2}}{\eta^{2}C^{2}}\right\}\right).$ (48)
@@
$\mathcal{D}_{\alpha}^{(z_{t+1})}(\mathbb{P}_{\boldsymbol{\theta}_{t+1}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{t+1}})$ (49)
$\leq\mathcal{D}^{(z_{t})}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{t}}||\mathbb{P}_{\boldsymbol{\theta}_{t}^{\prime}})+\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}+\frac{\alpha a_{t}^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}.$
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})$ (50)
$\leq\mathcal{D}_{\alpha}^{(z_{\tau})}(\mathbb{P}_{\boldsymbol{\theta}_{\tau}}||\mathbb{P}_{\boldsymbol{\theta}_{\tau}^{\prime}})+\sum_{t=\tau}^{T-1}\frac{\alpha a_{t}^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}+\sum_{t=\tau}^{T-1}\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}$
@@
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})$ (51)
$\leq\min_{\tau\in\{0,\cdots,T-1\}}\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}(T-\tau)+\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}$
@@
Define
$g(\beta)=\frac{2\alpha C^{2}}{\beta nb\sigma_{\text{DP}}^{2}}+\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}(1-\beta)}.$ (52)
By setting $g^{\prime}(\beta)=0$, one may obtain
$\beta^{*}=\frac{\sqrt{\frac{2\alpha C^{2}}{nb\sigma_{\text{DP}}^{2}}}}{\sqrt{\frac{2\alpha C^{2}}{nb\sigma_{\text{DP}}^{2}}}+\sqrt{\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}}}}.$ (53)
Substituting $\beta^{*}$ into the original optimization problems, it becomes
$g(\beta^{*})=\left(\sqrt{\frac{2\alpha C^{2}}{nb\sigma_{\text{DP}}^{2}}}+\sqrt{\frac{\alpha(1+\eta L)^{2}D^{2}}{2\eta^{2}\sigma_{\text{DP}}^{2}}}\right)^{2}.$ (54)
Combined with the linear bound in (46), we can obtain
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}_{T}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime}_{T}})\!\leq\!\mathcal{O}\left(\frac{\alpha C^{2}}{nb\sigma_{\textup{DP}}^{2}}\!\min\left\{T,\frac{(1+\eta L)^{2}nbD^{2}}{2\eta^{2}C^{2}}\right\}\right).$ (55)
@@ Appendix C Proofs for Utility Analysis
Our utility analysis builds upon the results of bassily2014private, zhang2020improved, and koloskova2023revisiting.
@@
where
$g(\boldsymbol{\theta}_{t})=\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}g_{\xi}(\boldsymbol{\theta}_{t})=\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}\operatorname{clip}_{C}\left(\nabla l_{\xi}(\boldsymbol{\theta}_{t})\right),$ (57)
and $\boldsymbol{\zeta}_{t}\sim\mathcal{N}(\boldsymbol{0},\eta^{2}\sigma_{\text{DP}}^{2}\boldsymbol{I}_{d})$ is the Gaussian perturbation.
@@
$\mathbb{E}_{\boldsymbol{\zeta}_{t}}\left[\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}^{*}\|^{2}\right]$ (58)
$\leq\mathbb{E}_{\boldsymbol{\zeta}_{t}}\left[\|\boldsymbol{\theta}_{t}-\eta g(\boldsymbol{\theta}_{t})+\boldsymbol{\zeta}_{t}-\boldsymbol{\theta}^{*}\|^{2}\right]$
@@
$\leq\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|^{2}-\frac{1}{b}\sum_{\xi\in\mathcal{B}_{t}}2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$+\eta^{2}C^{2}+d\eta^{2}\sigma_{\text{DP}}^{2},$
@@
$-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$ (59)
$=-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma_{\xi}\nabla l(\boldsymbol{\theta}_{t})+\gamma_{\xi}\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$\leq-2\eta\gamma_{\xi}[\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})-\gamma_{\xi}\nabla l(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$
$\leq-\frac{7\eta C}{4\|\nabla l(\boldsymbol{\theta}_{t})\|}[l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})]$
$+2\eta\gamma_{\xi}\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|\cdot\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|$
$\leq-\frac{7\eta C}{4\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
$+\frac{2\eta C\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{4}}}{7\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{3}{4}}-\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{4}}}\sqrt{\frac{2}{\mu}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
$\leq-\frac{7\eta C}{4\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
$+\frac{4\eta C}{\sqrt{2}(7\sqrt{L}-\sqrt{\mu})}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}$
@@
where the third step is by the convexity and Cauchy–Schwarz inequality, and the fourth step is by Lemma A.14 and A.15.
@@
$-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})$ $\leq 2\eta C\|\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*}\|$ (60)
$\leq\frac{4\eta C}{\sqrt{2\mu}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})},$
by using Cauchy–Schwarz inequality and Lemma A.15. Here, we define $\kappa_{\xi}=\boldsymbol{1}\{\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|>5\sigma_{\text{SGD}}(\frac{L}{\mu})^{\frac{1}{4}}\}$. According to Lemma A.2, we have
$\operatorname{Pr}(\kappa_{\xi}=1)$ $\!=\!\operatorname{Pr}\left(\|\nabla l_{\xi}(\boldsymbol{\theta}_{t})-\nabla l(\boldsymbol{\theta}_{t})\|^{2}>25\sigma_{\text{SGD}}^{2}\sqrt{L/\mu}\right)$ (61)
$\leq\frac{\sigma_{\text{SGD}}^{2}}{25\sigma_{\text{SGD}}^{2}\sqrt{L/\mu}}=\frac{1}{25}\sqrt{\frac{\mu}{L}},$
@@
$\mathbb{E}_{\xi}\left[-2\eta[g_{\xi}(\boldsymbol{\theta}_{t})]^{T}(\boldsymbol{\theta}_{t}-\boldsymbol{\theta}^{*})\right]$ (62)
$\leq-\frac{13\eta C}{12\sqrt{2L}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\cdot\frac{24}{25}$
$+\frac{4\eta C}{\sqrt{2\mu}}\sqrt{l(\boldsymbol{\theta}_{t})-l(\boldsymbol{\theta}^{*})}\cdot\frac{1}{25}\sqrt{\frac{\mu}{L}}$